Please rank the stability of 1-butene, trans-2-butene, and cis-2-butene, and explain the reasoning behind the heats of hydrogenation, as it relates to each molecule.

In your own words, explain the difference between 1,2 and 1,4 additions. Please include terms of allylic carbocation, delocalization, and resonance. 3-5 sentences will be sufficient.

Please explain, detail, the differences in the hydrohalogenation across the double bond with regards to Markovnikov and Non-Markovnikov

this is all one question please type answer i cannot always understand handwriting
Answer #1

Stability order of alkenes :

trans-2-butene > cis-2-butene > 1- butene .

Stability of alkene is related to the number of C-H bonds. More the C-H bonds more would be hyperconjugation and more stability of alkene. Trans alkene is more stable than cis alkene due to steric factor arises in cis alkene.

Heat of hydrogenation is inversely proportional to the stability of alkene.

CH2=CH-CH=CH2 + HBr CH2=CH-CH(Br)CH3 (1,2 product) + CH2(Br)CH=CH-CH3(1,4 product )

1,2 addition is also called kinetic controlled product. In the mechanism allylic 2° carbocation will form which is more stable than the allylic 1° carbocation. Allylic carbocation is resonance stabilized. Activation energy would be less for 1,2 addition , hence it would faster compared to 1,4 addition. 1,4 addition product is thermodynamically more stable than 1,2 addition. At higher temperature 1,4 addition would be major product.

When HX is added to an unsymmetrical alkene, the X- would attach to that carbon , which has lesser number of hydrogen atom. The reaction is electrophilic addition reaction. First step is formation of carbocation. Stability of carbocation is 3°>2°>1°. Then in second step nucleophile (X-) would attack on carbocation.

While addition of HBr in presence of peroxide, to alkene would be opposite of markovnikov rule. Antimarkovnikov rule.

CH3-CH=CH2 + HBr +( ROOR) CH3CH2CH2Br